A passenger train travels 295 mi in the same amount of time it takes a freight train to travel 225 mi. The rate of the passenger train is 14 mph greater than the rate of the freigh train. Find the rate of each train.
Distance Rate Time
passenger 295 14+r 295
14+r


freight 295 r 225 225
r

295 + 225 = ?
14+r r

Answer by edjones(8007)

Let f=rate of the freight train, r=rate, d=distance, t=time
rt=d
t=d/r
295/(f+14)=225/f Since the times are equal.
295f=225(f+14)
295f=225f+3150
70f=3150
f=45 mph
f+14=59 mph
.
Ed